OpenStudy (anonymous):

Solve a = 2r - 6p for p.

OpenStudy (anonymous):

first isolate the p term so its alone on one side then solve for p

OpenStudy (jhannybean):

\[\large a = 2r - 6p \]subtract 2r from both sides. \[\large a-2r = -2r + 2r -6p\] simplify. \[\large a-2r = -6p\] Now divide 6 to both sides. \[\large \frac{a-2r}{-6}= p\]Simplify the numerator by dividing each term by -6.\[\large -\frac{a}{6}+\frac{r}{3}=p\]